OpenStudy (anonymous):

find how many quarts of 4% butterfat milk and 1 % butterfat milk should be mixed to yield 75% quarts of 2 % butterfat milk. i have no idea how to even begin this problem...

OpenStudy (mertsj):

You need to state the problem correctly. 75 % quarts of 2% butterfat milk makes no sense.

OpenStudy (anonymous):

i meant 75 quarts not %...sorry...

OpenStudy (mertsj):

Let x = number of quarts of 4% milk to use. Then 75-x = number of quarts of 1% milk to use. .04x = butterfat in the x quarts of 4% milk .01(75-x)= butterfat in the 75-x quarts of 1% milk .02(75)= butterfat in the final 75 quarts of milk. So: .04x+.01(74-x)=.02(75) Solve it and you will have the number of quarts of 4% milk to use.

OpenStudy (anonymous):

why is it 74-x?

OpenStudy (mertsj):

It's not. It's 75-x

OpenStudy (mertsj):

Sorry for the typo

OpenStudy (anonymous):

oh ok! its fine

OpenStudy (anonymous):

thank you!!!!!!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

do you do distributive property first?

OpenStudy (mertsj):

I would multiply each term by 100 to get rid of those decimals

OpenStudy (anonymous):

event he numbers that aren't decimals multiply by 100?

OpenStudy (anonymous):

including x?

OpenStudy (mertsj):

each term. There are two terms on the left and one term on the right

OpenStudy (anonymous):

i dont understand

OpenStudy (mertsj):